# CHIPS-FF: Evaluating Universal Machine Learning Force Fields for Material Properties

### v3

@@ Abstract
Multiscale modeling in materials science 1, 2 is an essential tool to bridge the gap between atomistic properties, device performance, and manufacturing. With the passage of the U.S. Creating Helpful Incentives to Produce Semiconductors (CHIPS) and Science Act in 2022 3, there has been a renewed focus on modeling semiconductor materials, devices, and components. Various approaches to semiconductor modeling can be applied at different scales, including quantum mechanical tools such as density functional theory (DFT)4, classical atomistic simulations such as molecular dynamics (MD) 5, technology computer-aided design (TCAD) 6 and various machine learning (ML) models 7, 8, 9, 10, 11, 12, 13.
@@
In order to better inform semiconductor device models, accurate structure-to-property relationships must be established, which often require more computationally expensive quantum mechanical approaches. Accurate structural relaxations of defect structures and semiconductor interfaces are key to uncovering the underlying physics that govern device performance. For example, defects in semiconductors can introduce energy levels within the band gap, which can impact carrier mobility and concentration and create traps, leading to non-radiative recombination of holes and electrons 14, 15, 16, 17, 18, 19. Semiconductor surfaces and interfaces are especially important to consider since they are the building blocks of devices, impacting charge carrier behavior (band alignment), electrical properties (i.e., Schottky vs. Ohmic), thermal management and device efficiency 20, 21, 22, 23, 24, 25. In addition, the combination of semiconductor materials and high-dielectric amorphous structures are essential for building Metal-Oxide-Semiconductor Field-Effect Transistors (MOSFETs) 26, 27, 28, 29, 30, 31.
@@
Due to the costly scaling of DFT (, where is the number of electrons), high-throughput calculations of larger simulation cells required to study defects, surfaces and interfaces are not feasible. Unfortunately, many of the current classical atomistic interatomic potentials do not have the capability to achieve near-DFT accuracy and capture these complex interactions. In addition, most classical potentials are limited in their transferability to various systems and experimental conditions. Early attempts to create a universal force field for the entire periodic table were limited in their accuracy 32. For this reason, many researchers have explored machine learning force fields (MLFFs) trained on highly-accurate DFT calculations as a viable route to bridge the gap between quantum mechanical accuracy and large-scale atomistic simulations.
@@
One of the pioneering MLFF architectures was developed by Behler and Parrinello in 2007 using neural networks 33. Other MLFF methods include Gaussian processes-based Gaussian approximation potentials (GAP) 34, spectral neighbor analysis potential (SNAP) 35 and Allegro 36. Graph neural network (GNN)-based machine learning force fields (MLFFs) are now considered state-of-the-art due to their superior accuracy and transferability. With the public availability of large DFT databases (such as the Materials Project 37, JARVIS-DFT 38, 39, OQMD 40, 41, and Alexandria 42, 43, 44, 45), several researchers have trained GNN force fields on full DFT databases, encompassing the entire periodic table. These so-called universal/unified/foundational machine learning force fields (uMLFFs) have been successful at reproducing near-DFT accuracy for a wide-range of systems.
@@
A major advantage of utilizing these pretrained uMLFF architectures is that expensive DFT dataset generation is not an initial requirement. These pretrained uMLFFs can be used in place of DFT or fine tuned on more tailored datasets or tasks. Such uMLFFs have shown remarkable potential for numerous applications including modeling of defects and surfaces 46, 47, disordered alloys 48, high-pressure superconductors 49, solid-solution energetics and ion migration barriers 46, catalysis 50, thermal conductivity 51, semiconductor interfaces 21, guiding inverse materials design 12, 52, 53 and many other applications in chemistry, physics and materials science 54, 55, 56. Important quantities such as optimized structure, elastic properties, vibrational properties (phonons), stability, thermal properties, defect formation energy and surface energy can be easily computed from uMLFFs at a significantly lower cost than DFT. Additionally, uMLFF structural relaxations of bulk semiconductors, point defects, surfaces and interfaces can be coupled with other low-cost tools (direct prediction with GNNs 57, 17, 21, 17, 58, tight binding 59, 60) to predict electronic properties. uMLFFs can also be used to perform structure searches (replacing hundreds to tens of thousands of DFT relaxation calculations), which can be followed by a single DFT calculation for electronic properties of the target structure.
@@
One of the first graph-based uMLFF architectures introduced in 2022 was M3GNet from Chen and Ong 61. M3GNet has since transformed into MatGL (Materials Graph Library) 62, a re-implementation of M3GNet built on the Deep Graph Library (DGL) 63 and PyTorch 64. M3GNet evolved from the MEGNet (MatErials Graph Network) property prediction model introduced in Ref. 65. M3GNet was trained on 60,000 inorganic materials ( 190,000 relaxation steps) in the Materials Project (MP), covering 89 elements of the periodic table 61. This dataset is commonly referred to as MPF 61. In M3GNet, the angles (representing three-body interactions) are incorporated by aggregating to bonded atoms within the graph convolution steps to update atoms, bonds and properties 61. The Atomistic LIne Graph Neural Network (ALIGNN)-FF model 66 was introduced in 2022 and was trained on the JARVIS-DFT dataset of 75,000 materials ( 300,000 relaxation steps, ALIGNN-FF DB). ALIGNN-FF evolved from the ALIGNN property prediction model released in 2021 58. The ALIGNN model achieved significant improvement in property prediction tasks due to the inclusion of bond-angles in terms of line graphs, which previous GNN models lacked 58. The earlier versions of ALIGNN-FF utilized a k-nearest neighbor graph 66, while the updated version of ALIGNN-FF (2024.12.2) utilizes a radius graph, making it faster and more robust. Another uMLFF that utilizes a crystal graph (atom graph and bond graph) is the Crystal Hamiltonian Graph Neural Network (CHGNet) 67, which was proposed in 2023. This model was trained on the Materials Project Trajectory (MPtrj) dataset 67, which includes DFT calculations for 150,000 materials ( 1.58 million relaxation steps). CHGNet also incorporates magnetic moments into the training process to enhance the description of chemical reactions 67.
@@
MACE68, 69, 54, an equivariant message passing neural network potential, uses higher order messages (higher than two body) combined with atomic cluster expansion (ACE) 70, 71, which is a method for deriving an efficient body-ordered symmetric polynomial basis to represent functions of atomic neighborhoods. A pre-trained model for the entire periodic table (trained on MPTrj) for MACE was released in late 2023 54 (referred to as MACE-MP-0). In late 2024, a new pretrained version of MACE (known as MACE-MPA-0) was released, which is trained on a combination of MPTrj and Alexandria 72. SevenNet (Scalable EquiVariance Enabled Neural Network) 73, another equivariant message passing uMLFF, is based on the NequIP 74 architecture for building E(3)-equivariant force fields. SevenNet trained on the MPtrj dataset and was released in 2024. In addition to these open source models, there have been proprietary uMLFF models developed including MatterSim (Microsoft) 75 and GNoME (Google DeepMind) 76. A version of MatterSim (MatterSim-v1), which has a similar but slightly modified architecture to M3GNet and is trained on over 17 million structures sampled from Materials Project, Alexandria, and Microsoft’s own dataset (consisting of additional structures and MD trajectories at ambient and extreme conditions), was made open source to the community in December 2024 77. The Orb 78 pretrained uMLFF, released in fall 2024, was developed by the startup company Orbital Materials, where superior performance for force and energy predictions were demonstrated at significantly reduced computational cost 55, 79, 80. Orb utilizes an attention augmented Graph Network-based Simulator (GNS) 81, which is a type of Message Passing Neural Network (MPNN) 82. In contrast to equivariant models such as MACE and SevenNet, Orb does not utilize equivariant message passing. The open source Orb model (orb-v2) is trained on MPtrj and Alexandria (over 3 million materials and 32 million relaxation steps) 55. In late 2024, Meta FAIR Chemistry released the Open Materials 2024 (OMat24) 83 models and dataset, which contains over 110 million DFT calculations and pretrained uMLFF models based on the EquiformerV2 84 architecture, which is a highly accurate equivariant transformer model. OMat24 offers various sized models trained on OMat, MPTrj and Alexandria.
@@
Recently, interest in the field of uMLFFs has grown immensely (from academia, government and industry), as seen by the several models above and various use cases of new model architectures. Given the current state of the field, benchmarking the performance of uMLLFs is imperative. Large-scale efforts to benchmark the performance of various pretrained uMLFF architectures have emerged such as Matbench Discovery 85, 80 and the JARVIS-Leaderboard 86, 87, which calculate error metrics and allow for new models to be uploaded as time progresses. There have been focused benchmarking efforts for more involved properties beyond energy. Yu et al. 88 performed a systematic assessment of uMLFFs (M3GNet, ALIGNN-FF, CHGNet and MACE), calculating properties such as optimized structure, formation energy, bulk modulus, and phonon band structure, comparing directly to DFT to calculate error metrics. Benchmarking of more advanced properties such as surface energies, defect formation energies, solid-solution energetics, and ion migration barriers have recently been performed with M3GNet, CHGNet and MACE 46, 47. In addition, a high-level assessment of uMLFF computed thermal conductivity has been conducted with M3GNet, CHGNet, MACE, SevenNet and ORB by Pot́a et al. 51 More recently, Zhu et al. utilized M3GNet, CHGNet, MACE, SevenNet, and ORB to accelerate the prediction of CALPHAD-based phase diagrams for complex alloys 89 and Lowe et al. performed in-depth benchmarking of phonons for over 10,000 materials using all of the recent state-of-the-art uMLFFs 90. These recent efforts have highlighted the successes and limitations of various uMLFF architectures, and have emphasized the need for a comprehensive publicly available benchmarking platform. Such a platform allows the community to make informed decisions on regarding uMLFFs, outweighing factors such as computational cost versus accuracy. In addition, researchers may find systematic trends in the uMLFF results which can help others to understand the applicability of each uMLFF type. Well-understood systematic trends in DFT functionals (i.e., the underbinding and lattice constant errors of GGA) have been exploited to identify new exfoliable van der Waals (vdW) materials 91. In contrast to the Jacob’s Ladder metaphor of DFT exchange-correlation functionals 92, where the “rungs” of the ladder refer to the increasing complexity (and computational cost) and accuracy of the functionals, there (as of yet) exists no systematic assessment or generalizable statements for uMLFFs. As uMLFFs begin to replace DFT calculations, such analysis is needed to help the community better develop new force fields or fine-tune existing force fields for downstream applications.
@@
In this work, we present CHIPS-FF (Computational High-performance Infrastructure for Predictive Simulation-based Force Fields): a generalized user-friendly open-source benchmarking package to test various uMLFFs for a number of properties beyond energy. We have implemented a streamlined workflow to use uMLFFs for structural relaxation, bulk modulus, elastic properties, point defect formation energy, surface energy, interfacial properties (work of adhesion), molecular dynamics and creation of amorphous structures, all with automated error metrics from JARVIS-DFT. In contrast to large scale benchmarking efforts such as Matbench Discovery 85, 80, which requires users to submit contributions covering enormous test sets (on the order of 200,000), CHIPS-FF allows for robust benchmarking on smaller datasets and more complex properties. Although our framework is agnostic to the type of material, we chose a set of 104 materials that are most relevant for semiconductor devices as a case study. Additionally, we evaluate force-prediction accuracy of these models for close to 2 million atomic structures from Materials Project and JARVIS-DFT databases. We intend this codebase and benchmarking dataset to benefit the MLFF community and aid in the development of future models.
@@
The CHIPS-FF workflow mainly utilizes the Atomic Simulation Environment (ASE) 93 and JARVIS-Tools 38, 39 to perform atomistic simulations in Python. In terms of pretrained uMLFF models, we utilized ALIGNN-FF: 2.12.2024 (trained on 307k entries from JARVIS-DFT), CHGNet: 0.3.8, , MatGL: 1.1.2 (using both the M3GNet-MP-2021.2.8-PES and M3GNet-MP-2021.2.8-DIRECT-PES 94 models), MACE: 0.3.10 (using MACE-MP-0 and MACE-MPA-0), SevenNet-0: 0.9.2 (using the 11July2024 model), orb-models: 0.4.1 (orb-v2 and orb-d3-v2, which utilizes the D3 dispersion correction 95), and MatterSim-v1 (MatterSim-v1.0.0-5M). In addition, we benchmarked a MACE force field trained on the newly released Alexandria dataset presented in Ref. 45 (this model was trained on over 200,000 2D material calculations with the PBEsol 96 functional and is referred to as MACE in that work 97). In our work, we refer to this model as mace-alexandria (which is distinct from mace-mpa). For the OMat24 models from Meta, we tested the three EquiformerV2 (eqV2) models trained on OMat (Small: 31M (million), medium: 86M, and large: 153M in terms of model parameters) and two eqV2 models trained on OMat+MPTrj+Alexandria (Small: 31M and medium: 86M) 56, 83. These models are named as: eqV2_31M_omat, eqV2_86M_omat, eqV2_153M_omat, eqV2_31M_omat_mp_salex, and eqV2_86M_omat_mp_salex 56, 83.
@@
Initial structures were obtained from the JARVIS-DFT database (but not limited by it) and then optimized using the FIRE 98 algorithm within ASE. For bulk relaxations, atomic coordinates and lattice vectors were allowed to relax, but for surfaces and vacancy calculations, cell volume was fixed and atomic positions were allowed to relax. The maximum force value (stopping criterion for relaxation) was set to 0.05 eV/Å. If a calculation did not reach the stopping criterion by 200 steps, it was considered unconverged. If a structure did not reach convergence within 200 steps, the final structure and energy at 200 steps was logged and used for subsequent portions of the workflow. The bulk modulus was obtained by applying isotropic strain to the system from -6 to +6 and using a Murnaghan fit for the equation of state. The elastic tensor was computed using the Elastic package 99, 100, 101. For bulk modulus and elastic tensor calculations, the conventional cell was used. We utilized the phonopy 102, 103 package to perform simulations of the phonon spectrum using the finite displacement method 104, where a 2 x 2 x 2 supercell was constructed from the initial unit cell and a displacement of 0.2 Å was used. From these phonopy results, we also extracted the Helmholtz free energy (and zero point energy at 0 K), entropy and heat capacity as a function of temperature. We have added capabilities to CHIPS-FF to compute the thermal expansion of each material at constant pressure and the lattice thermal conductivity, utilizing the quasi-harmonic approximation 105 implemented in phonopy and phono3py 102, 103 respectively (these simulations are ongoing and will be discussed in future work). JARVIS-Tools was used to automatically generate supercells for neutral point defects (vacancies) and non-polar surfaces (both of which were later relaxed with uMLFFs). Defect supercells are constructed from the conventional cell and enforced to have a lattice vector c as close to 8 Å as possible, where single vacancies are created based on Wycoff-position information for each atomic species in the crystal. The following formula is used to obtain the vacancy formation energy ():
@@
where is the energy of the surface structure, is the energy of the bulk structure, is the number of bulk unit cells in the surface, and is the cross sectional area of the surface. In order to study amorphous materials, we have implemented finite-temperature molecular dynamics into our workflow to perform melt/quench simulations using Berendsen NVT (Number of particles, Volume, Temperature ensemble) dynamics within ASE. For our case study of amorphous Si, we used a timestep of 1 femtosecond and ran the simulation at 3500 K for 10 ps (melt) and 300 K for 20 ps (quench). For benchmarking calculations to compare uMLFF results to, we performed similar ab initio MD (AIMD) calculations with the Vienna Ab initio Simulation Package (VASP) code, using projector augmented wave (PAW) pseudopotentials 106, 107 and the vdW-DF-optB88 functional. In order to perform calculations for material interfaces, we utilized InterMat 21, a newly developed package for the generation and calculation of interface structures (substrate plus film). The initial interface is obtained by creating a superlattice or alternating slab junction (ASJ) structure (without vacuum padding) 108, 109 and then using the Zur algorithm 110 to obtain the best candidate interface. From here, we computed the relative alignment between the film and substrate in the in-plane (xy) direction by performing a grid search with a 0.05 fractional spacing interval with each respective uMLFF model. In addition to determining the optimal in-plane orientation between the substrate and film, this gives us a qualitative estimate of how smooth the potential energy surface is at the interface for each pretrained uMLFF. From this calculation, we also determined the work of adhesion () at the interface with:

### v4

@@ Abstract
Multiscale modeling in materials science 1, 2 is an essential tool to bridge the gap between atomistic properties, device performance, and manufacturing. Recently, there has been a renewed focus on modeling semiconductor materials, devices, and components. Various approaches to semiconductor modeling can be applied at different scales, including quantum mechanical tools such as density functional theory (DFT)3, classical atomistic simulations such as molecular dynamics (MD) 4, technology computer-aided design (TCAD) 5 and various machine learning (ML) models 6, 7, 8, 9, 10, 11, 12.
@@
In order to better inform semiconductor device models, accurate structure-to-property relationships must be established, which often require more computationally expensive quantum mechanical approaches. Accurate structural relaxations of defect structures and semiconductor interfaces are key to uncovering the underlying physics that govern device performance. For example, defects in semiconductors can introduce energy levels within the band gap, which can impact carrier mobility and concentration and create traps, leading to non-radiative recombination of holes and electrons 13, 14, 15, 16, 17, 18. Semiconductor surfaces and interfaces are especially important to consider since they are the building blocks of devices, impacting charge carrier behavior (band alignment), electrical properties (i.e., Schottky vs. Ohmic), thermal management and device efficiency 19, 20, 21, 22, 23, 24. In addition, the combination of semiconductor materials and high-dielectric amorphous structures are essential for building Metal-Oxide-Semiconductor Field-Effect Transistors (MOSFETs) 25, 26, 27, 28, 29, 30.
@@
Due to the costly scaling of DFT (, where is the number of electrons), high-throughput calculations of larger simulation cells required to study defects, surfaces and interfaces are not feasible. Unfortunately, many of the current classical atomistic interatomic potentials do not have the capability to achieve near-DFT accuracy and capture these complex interactions. In addition, most classical potentials are limited in their transferability to various systems and experimental conditions. Early attempts to create a universal force field for the entire periodic table were limited in their accuracy 31. For this reason, many researchers have explored machine learning force fields (MLFFs) trained on highly-accurate DFT calculations as a viable route to bridge the gap between quantum mechanical accuracy and large-scale atomistic simulations.
@@
One of the pioneering MLFF architectures was developed by Behler and Parrinello in 2007 using neural networks 32. Other MLFF methods include Gaussian processes-based Gaussian approximation potentials (GAP) 33, spectral neighbor analysis potential (SNAP) 34 and Allegro 35. Graph neural network (GNN)-based machine learning force fields (MLFFs) are now considered state-of-the-art due to their superior accuracy and transferability. With the public availability of large DFT databases (such as the Materials Project 36, JARVIS-DFT 37, 38, OQMD 39, 40, and Alexandria 41, 42, 43, 44), several researchers have trained GNN force fields on full DFT databases, encompassing the entire periodic table. These so-called universal/unified/foundational machine learning force fields (uMLFFs) have been successful at reproducing near-DFT accuracy for a wide-range of systems.
@@
A major advantage of utilizing these pretrained uMLFF architectures is that expensive DFT dataset generation is not an initial requirement. These pretrained uMLFFs can be used in place of DFT or fine tuned on more tailored datasets or tasks. Such uMLFFs have shown remarkable potential for numerous applications including modeling of defects and surfaces 45, 46, disordered alloys 47, high-pressure superconductors 48, solid-solution energetics and ion migration barriers 45, catalysis 49, thermal conductivity 50, semiconductor interfaces 20, guiding inverse materials design 11, 51, 52 and many other applications in chemistry, physics and materials science 53, 54, 55. Important quantities such as optimized structure, elastic properties, vibrational properties (phonons), stability, thermal properties, defect formation energy and surface energy can be easily computed from uMLFFs at a significantly lower cost than DFT. Additionally, uMLFF structural relaxations of bulk semiconductors, point defects, surfaces and interfaces can be coupled with other low-cost tools (direct prediction with GNNs 56, 16, 20, 16, 57, tight binding 58, 59) to predict electronic properties. uMLFFs can also be used to perform structure searches (replacing hundreds to tens of thousands of DFT relaxation calculations), which can be followed by a single DFT calculation for electronic properties of the target structure.
@@
One of the first graph-based uMLFF architectures introduced in 2022 was M3GNet from Chen and Ong 60. M3GNet has since transformed into MatGL (Materials Graph Library) 61, a re-implementation of M3GNet built on the Deep Graph Library (DGL) 62 and PyTorch 63. M3GNet evolved from the MEGNet (MatErials Graph Network) property prediction model introduced in Ref. 64. M3GNet was trained on 60,000 inorganic materials ( 190,000 relaxation steps) in the Materials Project (MP), covering 89 elements of the periodic table 60. This dataset is commonly referred to as MPF 60. In M3GNet, the angles (representing three-body interactions) are incorporated by aggregating to bonded atoms within the graph convolution steps to update atoms, bonds and properties 60. The Atomistic LIne Graph Neural Network (ALIGNN)-FF model 65 was introduced in 2022 and was trained on the JARVIS-DFT dataset of 75,000 materials ( 300,000 relaxation steps, ALIGNN-FF DB). ALIGNN-FF evolved from the ALIGNN property prediction model released in 2021 57. The ALIGNN model achieved significant improvement in property prediction tasks due to the inclusion of bond-angles in terms of line graphs, which previous GNN models lacked 57. The earlier versions of ALIGNN-FF utilized a k-nearest neighbor graph 65, while the updated version of ALIGNN-FF (2024.12.2) utilizes a radius graph, making it faster and more robust. Another uMLFF that utilizes a crystal graph (atom graph and bond graph) is the Crystal Hamiltonian Graph Neural Network (CHGNet) 66, which was proposed in 2023. This model was trained on the Materials Project Trajectory (MPtrj) dataset 66, which includes DFT calculations for 150,000 materials ( 1.58 million relaxation steps). CHGNet also incorporates magnetic moments into the training process to enhance the description of chemical reactions 66.
@@
MACE67, 68, 53, an equivariant message passing neural network potential, uses higher order messages (higher than two body) combined with atomic cluster expansion (ACE) 69, 70, which is a method for deriving an efficient body-ordered symmetric polynomial basis to represent functions of atomic neighborhoods. A pre-trained model for the entire periodic table (trained on MPTrj) for MACE was released in late 2023 53 (referred to as MACE-MP-0). In late 2024, a new pretrained version of MACE (known as MACE-MPA-0) was released, which is trained on a combination of MPTrj and Alexandria 71. SevenNet (Scalable EquiVariance Enabled Neural Network) 72, another equivariant message passing uMLFF, is based on the NequIP 73 architecture for building E(3)-equivariant force fields. SevenNet trained on the MPtrj dataset and was released in 2024. In addition to these open source models, there have been proprietary uMLFF models developed including MatterSim (Microsoft) 74, GNoME (Google DeepMind) 75, and PFP (Preferred Networks Inc) 76. A version of MatterSim (MatterSim-v1), which has a similar but slightly modified architecture to M3GNet and is trained on over 17 million structures sampled from Materials Project, Alexandria, and Microsoft’s own dataset (consisting of additional structures and MD trajectories at ambient and extreme conditions), was made open source to the community in December 2024 77. The Orb 78 pretrained uMLFF, released in fall 2024, was developed by the startup company Orbital Materials, where superior performance for force and energy predictions were demonstrated at significantly reduced computational cost 54, 79, 80. Orb utilizes an attention augmented Graph Network-based Simulator (GNS) 81, which is a type of Message Passing Neural Network (MPNN) 82. In contrast to equivariant models such as MACE and SevenNet, Orb does not utilize equivariant message passing. The open source Orb model (orb-v2) is trained on MPtrj and Alexandria (over 3 million materials and 32 million relaxation steps) 54. In late 2024, Meta FAIR Chemistry released the Open Materials 2024 (OMat24) 83 models and dataset, which contains over 110 million DFT calculations and pretrained uMLFF models based on the EquiformerV2 84 architecture, which is a highly accurate equivariant transformer model. OMat24 offers various sized models trained on OMat, MPTrj and Alexandria.
@@
Recently, interest in the field of uMLFFs has grown immensely (from academia, government and industry), as seen by the several models above and various use cases of new model architectures. Given the current state of the field, benchmarking the performance of uMLLFs is imperative. Large-scale efforts to benchmark the performance of various pretrained uMLFF architectures have emerged such as Matbench Discovery 85, 80 and the JARVIS-Leaderboard 86, 87, which calculate error metrics and allow for new models to be uploaded as time progresses. Matbench Discovery is an interactive leaderboard platform where uMLFFs are ranked based on accuracy of properties (such as energy and more recently thermal conductivity) and the ability to simulate high-throughput discovery of stable new inorganic materials. These uMLFFs are evaluated on targeted large-scale test datasets (i.e., a test set of 200,000 Materials Project entries) to calculate error metrics 85, 80. The JARVIS-Leaderboard is a flexible benchmarking platform that allows for the evaluation of various machine learning, electronic structure, force-field and quantum computing, and experimental methods. So far, there are over 300 benchmarks and over 2,000 user contributions 86, 87. A subset of these benchmarks and user contributions are focused on uMLFFs for various test datasets 86, 87. There have been focused benchmarking efforts for more involved properties beyond energy. Yu et al. 88 performed a systematic assessment of uMLFFs (M3GNet, ALIGNN-FF, CHGNet and MACE), calculating properties such as optimized structure, formation energy, bulk modulus, and phonon band structure, comparing directly to DFT to calculate error metrics. Benchmarking of more advanced properties such as surface energies, defect formation energies, solid-solution energetics, and ion migration barriers have recently been performed with M3GNet, CHGNet and MACE 45, 46. In addition, a high-level assessment of uMLFF computed thermal conductivity has been conducted with M3GNet, CHGNet, MACE, SevenNet and ORB by Pot́a et al. 50 More recently, Zhu et al. utilized M3GNet, CHGNet, MACE, SevenNet, and ORB to accelerate the prediction of CALPHAD-based phase diagrams for complex alloys 89 and Lowe et al. performed in-depth benchmarking of phonons for over 10,000 materials using all of the recent state-of-the-art uMLFFs 90. These recent efforts have highlighted the successes and limitations of various uMLFF architectures, and have emphasized the need for a comprehensive publicly available benchmarking platform. Such a platform allows the community to make informed decisions on regarding uMLFFs, outweighing factors such as computational cost versus accuracy. In addition, researchers may find systematic trends in the uMLFF results which can help others to understand the applicability of each uMLFF type. Well-understood systematic trends in DFT functionals (i.e., the underbinding and lattice constant errors of GGA) have been exploited to identify new exfoliable van der Waals (vdW) materials 91. In contrast to the Jacob’s Ladder metaphor of DFT exchange-correlation functionals 92, where the “rungs” of the ladder refer to the increasing complexity (and computational cost) and accuracy of the functionals, there (as of yet) exists no systematic assessment or generalizable statements for uMLFFs. As uMLFFs begin to replace DFT calculations, such analysis is needed to help the community better develop new force fields or fine-tune existing force fields for downstream applications.
@@
In this work, we present CHIPS-FF (Computational High-performance Infrastructure for Predictive Simulation-based Force Fields): a generalized user-friendly open-source benchmarking package to test various uMLFFs for a number of properties beyond energy. We have implemented a streamlined workflow to use uMLFFs for structural relaxation, bulk modulus, elastic properties, point defect formation energy, surface energy, interfacial properties (work of adhesion), molecular dynamics and creation of amorphous structures, all with automated error metrics from JARVIS-DFT. In contrast to large scale benchmarking efforts such as Matbench Discovery 85, 80, which requires users to submit contributions covering large-scale test sets (on the order of 200,000), CHIPS-FF allows for robust benchmarking on smaller datasets and more complex properties. Although our framework is agnostic to the type of material, we chose a set of 104 materials that are most relevant for semiconductor devices as a case study. Additionally, we evaluate force-prediction accuracy of these models for close to 2 million atomic structures from Materials Project and JARVIS-DFT databases. We intend this codebase and benchmarking dataset to benefit the MLFF community and aid in the development of future models.
@@
The CHIPS-FF workflow mainly utilizes the Atomic Simulation Environment (ASE) 93 and JARVIS-Tools 37, 38 to perform atomistic simulations in Python. In terms of pretrained uMLFF models, we utilized ALIGNN-FF: 2.12.2024, CHGNet: 0.3.8, MatGL: 1.1.2 (using both the M3GNet-MP-2021.2.8-PES and M3GNet-MP-2021.2.8-DIRECT-PES 94 models), MACE: 0.3.10 (using MACE-MP-0 and MACE-MPA-0), SevenNet-0: 0.9.2 (using the 11July2024 model), orb-models: 0.4.1 (orb-v2 and orb-d3-v2, which utilizes the D3 dispersion correction 95), and MatterSim-v1 (MatterSim-v1.0.0-5M). In addition, we benchmarked a MACE force field trained on the newly released Alexandria dataset presented in Ref. 44 (this model was trained on over 200,000 2D material calculations with the PBEsol 96 functional and is referred to as MACE in that work 97). In our work, we refer to this model as mace-alexandria (which is distinct from mace-mpa). For the OMat24 models from Meta, we tested the three EquiformerV2 (eqV2) models trained on OMat (Small: 31M (million), medium: 86M, and large: 153M in terms of model parameters) and two eqV2 models trained on OMat+MPTrj+Alexandria (Small: 31M and medium: 86M) 55, 83. These models are named as: eqV2_31M_omat, eqV2_86M_omat, eqV2_153M_omat, eqV2_31M_omat_mp_salex, and eqV2_86M_omat_mp_salex 55, 83.
@@
Initial structures were obtained from the JARVIS-DFT database (but not limited by it) and then optimized using the FIRE 98 algorithm within ASE. For bulk relaxations, atomic coordinates and lattice vectors were allowed to relax, but for surfaces and vacancy calculations, cell volume was fixed and atomic positions were allowed to relax. The maximum force value (stopping criterion for relaxation) was set to 0.05 eV/Å. If a calculation did not reach the stopping criterion by 200 steps, it was considered unconverged. If a structure did not reach convergence within 200 steps, the final structure and energy at 200 steps was logged and used for subsequent portions of the workflow. We tested both the FrechetCellFilter (main text) and ExpCellFilter (in SI) within ASE 99. The bulk modulus was obtained by applying isotropic strain to the system from -6 to +6 and using a Murnaghan fit for the equation of state. The elastic tensor was computed using the Elastic package 100, 101, 102. For bulk modulus and elastic tensor calculations, the conventional cell was used. We utilized the phonopy 103, 104 package to perform simulations of the phonon spectrum using the finite displacement method 105, where a 2 x 2 x 2 supercell was constructed from the initial unit cell. Displacement values of 0.2 Å, 0.05 Å, 0.01 Å, and 0.001 Å were benchmarked. From these phonopy results, we also extracted the Helmholtz free energy (and zero point energy at 0 K), entropy and heat capacity as a function of temperature. We have added capabilities to CHIPS-FF to compute the thermal expansion of each material at constant pressure and the lattice thermal conductivity, utilizing the quasi-harmonic approximation 106 implemented in phonopy and phono3py 103, 104 respectively (these simulations are ongoing and will be discussed in future work). JARVIS-Tools was used to automatically generate supercells for neutral point defects (vacancies) and non-polar surfaces (both of which were later relaxed with uMLFFs). In this work, we obtained initial defect structures from the relaxed structures in the JARVIS-DFT vacancy database from Ref. 15. This allowed us to make a 1:1 comparison between vacancy formation energies (calculated with DFT and uMLFFs) for the exact same crystal structure and supercell size (avoiding finite-size scaling effects). In Ref. 15, these defect supercells were constructed from the conventional cell and enforced to have a lattice vector c of at least 8 Å , where single vacancies were created based on Wycoff-position information for each atomic species in the crystal. The following formula is used to obtain the vacancy formation energy ():
@@
where is the energy of the surface structure, is the energy of the bulk structure, is the number of bulk unit cells in the surface, and is the cross sectional area of the surface. In order to study amorphous materials, we have implemented finite-temperature molecular dynamics into our workflow to perform melt/quench simulations using Berendsen NVT (Number of particles, Volume, Temperature ensemble) dynamics within ASE. For our case study of amorphous Si, we used a timestep of 1 femtosecond and ran the simulation at 3500 K for 10 ps (melt) and 300 K for 20 ps (quench). For benchmarking calculations to compare uMLFF results to, we performed similar ab initio MD (AIMD) calculations with the Vienna Ab initio Simulation Package (VASP) code, using projector augmented wave (PAW) pseudopotentials 107, 108 and the vdW-DF-optB88 functional. These gamma-point AIMD calculation used a plane wave cutoff energy of 500 eV and the Nosé-Hoover109, 110, 111 thermostat to run NVT simulations at 2000 K for 5 ps (melting), and 300 K for 5 ps (at a timestep of 1 femtosecond). In order to perform calculations for material interfaces, we utilized InterMat 20, a newly developed package for the generation and calculation of interface structures (substrate plus film). The initial interface is obtained by creating a superlattice or alternating slab junction (ASJ) structure (without vacuum padding) 112, 113 and then using the Zur algorithm 114 to obtain the best candidate interface. From here, we computed the relative alignment between the film and substrate in the in-plane (xy) direction by performing a grid search with a 0.05 fractional spacing interval with each respective uMLFF model. In addition to determining the optimal in-plane orientation between the substrate and film, this gives us a qualitative estimate of how smooth the potential energy surface is at the interface for each pretrained uMLFF. From this calculation, we also determined the work of adhesion () at the interface with: